dgreer4:

Because carpeting typically comes in 12-ft widths the carpet for a 37ft by 16ft room must be laid out as shown to the right. a) Determine the total area of carpet needed for this room. B) at 6.45 per square foot installed, how much will it cost to carpet the room?

dgreer4:

1 attachment
dgreer4:

Here is the image please help if you can

Joe348:

To determine the total area of carpet needed for the room, we can multiply the length and width of the room. So, for a 37ft by 16ft room, the total area would be 37ft x 16ft = 592 square feet. To calculate the cost of carpeting the room at $6.45 per square foot installed, we can multiply the total area by the cost per square foot. So, the cost would be 592 square feet x $6.45 = $3,817.40. Therefore, it would cost approximately $3,817.40 to carpet the room. @dgreer4

surjithayer:

area of the room \[=37\frac{ 1 }{ 4 }\times 16=\frac{ 149 }{ 4}\times 16=596 ~sq.ft\] width of carpet =596 sq.ft. length of carpet \[=596\div 12=\frac{ 149 }{ 3 }=49 \frac{ 2 }{ 3 }~ft\] so he has to buy 50 ft long and 12 ft wide. or he needs to buy 600 sq. ft.

Spectrum:

Can you close this post? Clears out the posts so we know who has been helped and who needs help.